# Boundary Control and Calderón type Inverse Problems in Non-local heat equation

### v2

@@ 2. Density Results:
where $\epsilon>0$ is a fix quantity, $h\in\mathrm{L}^{2}(\Omega)$ is a given function and $\eta_{\tilde{T}}\in C_{c}^{\infty}((0,\tilde{T}))$ such that $\eta_{\tilde{T}}\equiv 1$ in $[T_{1},T_{2}]\subset(0,\tilde{T})$. Furthermore $u_{F^{*}}$ and $F^{*}$ are connected by (13) with $F^{*}$ as the initial data at $T=\tilde{T}$. We intend to show the functional ${\mathcal{J}_{\epsilon,a}:\mathrm{L}^{2}(\Omega)\to\mathbb{R}\cup\{+\infty\}}$ is lower semicontinuous, strictly convex and coercive. Just like the case of theorem 2 continuity of $\mathcal{J}_{\epsilon,a}$ follows from the eigenvalue representation and (8). Here (8) provides the boundary continuity. Strict convexity follows from the strict convexity of the $\mathrm{L}^{2}$-norm and observability estimate (theorem 9). We only prove coercivity of the functional $\mathcal{J}_{\epsilon,a}$.
@@
    Coercivity: Let $F_{m}^{*}\in\mathrm{L}^{2}(\Omega)$ for all $m\in\mathbb{N}$ and ${\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}\to+\infty}$ as $m\to+\infty$. Consider the functions ${G_{m}^{*}=\frac{F_{m}^{*}}{\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}$. Then ${\|G_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}=1}$ for all $m\in\mathbb{N}$. Let $u_{G^{*}_{m}}$ be the solution of (13) with respect to the initial data $G_{m}^{*}$. Then ${u_{G^{*}_{m}}=\frac{u_{F^{*}_{m}}}{\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}$. We divide the functional $\mathcal{J}_{\epsilon,a}$ by $\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}$. It yields
@@
    The second scenario is ${\liminf_{m\in\mathbb{N}}\int_{0}^{\tilde{T}}\int_{\partial\Omega}\eta_{\tilde{T}}^{2}\left|\frac{u_{G_{m}^{*}}}{d^{s}}\right|^{2}\,\rm{d}x\,\rm{d}t=0}$. Thanks to the observability estimate (9) we have that ${\|u_{G_{m}^{*}}\|_{\mathrm{L}^{2}((T_{1},T_{2})\times\Omega)}\to 0}$ as $m\to+\infty$ for all $T_{1}<T_{2}\in[0,\tilde{T}]$. We represents the solution $u_{G_{m}^{*}}$ as
@@
    Hence ${\langle\phi_{n},G_{m}^{*}\rangle\to 0}$ as $m\to+\infty$. As $G_{m}^{*}$ is uniformly bounded in $\mathrm{L}^{2}$, there exists $G^{*}\in\mathrm{L}^{2}(\Omega)$ such $G_{m}^{*}\rightharpoonup G^{*}$ weakly in $\mathrm{L}^{2}(\Omega)$. This further implies
@@
Continuity, strict convexity and coercivity helps us conclude that $\mathcal{J}_{\epsilon,a}$ attains it’s minimum for some $F_{min}^{*}\in\mathrm{L}^{2}(\Omega)$. We fix a function $G^{*}\in\mathrm{L}^{2}(\Omega)$ and let $u_{G^{*}}$ be the solution of (13) corresponding to the initial data $G^{*}$. We have
@@
$\frac{1}{2}$ $\int_{0}^{\tilde{T}}\int_{\Omega}2\alpha\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t+\frac{1}{2}\int_{0}^{\tilde{T}}\int_{\Omega}\alpha^{2}\eta_{\tilde{T}}^{2}\frac{(u_{G^{*}})^{2}}{d^{2a}}\,\rm{d}x\,\rm{d}t$
@@
$\int_{0}^{\tilde{T}}\int_{\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t+\epsilon\frac{\int_{\Omega}F^{*}_{min}G^{*}\,\rm{d}x}{\|F_{min}^{*}\|_{\mathrm{L}^{2}(\Omega)}}-\int_{\Omega}hG^{*}\,\rm{d}x\geq 0.$
Hence employingl Cauchy-Schwartz inequality, we obtain
$\int_{0}^{\tilde{T}}\int_{\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t-\int_{\Omega}hG^{*}\,\rm{d}x\geq-\epsilon\|G^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
@@
$\int_{0}^{\tilde{T}}\int_{\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t-\int_{\Omega}hG^{*}\,\rm{d}x\leq\epsilon\|G^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
@@
(29) $\left|\int_{0}^{\tilde{T}}\int_{\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t-\int_{\Omega}hG^{*}\,\rm{d}x\right|\leq\epsilon\|G^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
We choose ${F=\eta_{\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}}$ for the singular boundary data in (2), where $\eta_{\Omega}\in C_{c}^{\infty}(\partial\Omega)$ to be chosen later. Let ${F_{1}=(1-\eta_{\Omega})\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}}$. Let $u_{F},u_{F_{1}}$ be the solutions to (2) with singular boundary datum $F$ and $F_{1}$ respectively. Thanks to (9), we can conclude $F\in C_{c}^{\infty}((0,T)\times\Omega)$. [see [FRRO16], [Gru23]]. Integration by parts (16) yields that:
@@
$\int_{\Omega}u_{F_{1}}(\tilde{T},\cdot)G^{*}(\cdot)\,\rm{d}x=-\Gamma(a)\Gamma(a-1)\int_{0}^{\tilde{T}}\int_{\partial\Omega}\frac{u_{F_{1}}}{d^{a-1}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}s.$
@@
[…]
$\frac{u_{F^{*}_{min}}}{d^{s}}(t,z)=\sum\limits_{n\in\mathbb{N}}e^{-(\tilde{T}-t)\lambda_{n}}\langle\phi_{n},F^{*}_{min}\rangle\frac{\phi_{n}}{d^{s}}(z),\quad\forall z\in\partial\Omega.$
We use the fact that ${\left|\frac{\phi_{n}}{d^{s}}\right|\leq C_{\Omega}\lambda_{n}^{w}}$, where $w\in\mathbb{N}$ and $C_{\Omega}$ two positive constants depends on the domain, the dimension and the exponent ‘$a$’ only [see [AGCV23], [CGCV21], [Gru15b], [CGCV22], [FRRO16], [CKS10]]. It yields
$\left\|\frac{u_{F^{*}_{min}}}{d^{a}}\right\|_{\mathrm{L}^{\infty}([T_{3},T_{4}]\times\partial\Omega)}\leq$ $C_{\Omega}\|F^{*}_{min}\|_{\mathrm{L}^{2}(\Omega)}\sum\limits_{n\in\mathbb{N}}e^{-\lambda_{n}(\tilde{T}-T_{4})}\lambda_{n}^{w}$
$\leq$ $C_{\Omega}\|F^{*}_{min}\|_{\mathrm{L}^{2}(\Omega)}C_{\tilde{T},T_{4},w},$
where $C_{\tilde{T},T_{4},w}$ a positive constant depends only on $\tilde{T}-T_{4}$ and $w$. Let $u_{G^{*}}$ be the solution of (13) with initial data $G^{*}$. Integration by parts (16) yields:
$\int_{\Omega}u_{F_{1}}(\tilde{T},\cdot)G^{*}\,\rm{d}x=$ $-\Gamma(a)\Gamma(a-1)\int_{0}^{\tilde{T}}\int_{\partial\Omega}\frac{u_{F_{1}}}{d^{a-1}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}s$
$=$ $-\Gamma(a)\Gamma(a-1)\int_{0}^{\tilde{T}}\int_{\partial\Omega}(1-\eta_{\Omega})\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}s.$
Employing Hölder inequality, we obtain
$\int_{\Omega}u_{F_{1}}(\tilde{T},\cdot)G^{*}(\cdot)\,\rm{d}x\leq$ $\mathcal{Z}\text{Vol}(|\text{supp}(1-\eta_{\Omega})|)\left\|\frac{u_{F^{*}_{min}}}{d^{a}}\right\|_{\mathrm{L}^{\infty}([T_{3},T_{4}]\times\partial\Omega)}$
$\times$ $\left\|\frac{u_{G^{*}_{min}}}{d^{a}}\right\|_{\mathrm{L}^{\infty}([T_{3},T_{4}]\times\partial\Omega)}$
$\leq$ $\mathcal{Z}\text{Vol}(|\text{supp}(1-\eta_{\Omega})|)\|F^{*}_{min}\|_{\mathrm{L}^{2}(\Omega)}\|G^{*}\|_{\mathrm{L}^{2}(\Omega)},$
where ${\mathcal{Z}:=\tilde{T}\big{|}\Gamma(s)\Gamma(s-1)\big{|}C_{\Omega}^{2}C_{\tilde{T},T_{4},w}^{2}}$. Since $G^{*}$ is arbitrary we have that
$\|u_{F_{1}}(\tilde{T},\cdot)\|_{\mathrm{L}^{2}(\Omega)}\leq\mathcal{Z}\text{Vol}(\text{supp}|1-\eta_{\Omega}|)\|F^{*}_{min}\|_{\mathrm{L}^{2}(\Omega)}.$
If $F_{min}^{*}\equiv 0$, then ${\|u_{F_{1}}(\tilde{T},\cdot)\|_{\mathrm{L}^{2}(\Omega)}=0}$. Hence we are done. Otherwise, let $\eta_{\Omega}\in C_{c}^{\infty}(\partial\Omega)$ such that $\eta\equiv 1$ in $\tilde{\partial\Omega}$ compactly contained in $\partial\Omega$ such that
$\text{Vol}\left(\left|\partial\Omega\setminus\tilde{\partial\Omega}\right|\right)\leq\frac{\epsilon}{\mathcal{Z}\|F^{*}_{min}\|_{\mathrm{L}^{2}(\Omega)}}.$
Hence, ${\|u_{F_{1}}(\tilde{T},\cdot)\|_{\mathrm{L}^{2}(\Omega)}\leq\epsilon}$. Thus we conclude our proof. ∎
We next move onto the proofs of the Calderón type inverse problems for the non-local heat equations. We recover the potentials by analyzing the solution at a particular time slice for (1) and by analyzing the solution on the boundary for (2).
@@ 3. Calderón type Inverse Problems:
$-\partial_{t}w_{i}+(-\Delta)^{a}w_{i}+q_{i}w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\overline{\Omega}\\ \lim\limits_{\begin{subarray}{c}x\to y\\ y\in\partial\Omega\end{subarray}}\frac{w_{i}}{d^{a-1}}(y)=&F\qquad\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ w_{i}(T,\cdot)=&0\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $F\in C_{c}^{\infty}((0,T)\times\Omega)$. Since ${w_{i}\in\mathrm{L}^{2}\left((0,T);H^{(a-1)(2a)}(\overline{\Omega})\right)}$ for $i=1,2$ [see theorem 8, theorem 7], integration by parts (16) yields
@@
$\int_{\Omega}\partial_{t}u_{i}w_{i}+w_{i}\partial_{t}u_{i}=\int_{\Omega}\partial_{t}\left(u_{i}w_{i}\right)=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{w_{i}}{d^{a-1}}\frac{u_{i}}{d^{a}},\quad i=1,2.$
@@
$-\partial_{t}w_{i}+(-\Delta)^{a}w_{i}+q_{i}w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\overline{\Omega}\\ \lim\limits_{\begin{subarray}{c}x\to y\\ y\in\partial\Omega\end{subarray}}\frac{w_{i}}{d^{a-1}}(y)=&F\qquad\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ w_{i}(T,\cdot)=&0\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $F\in C_{c}^{\infty}((0,T)\times\Omega)$. Let
$w_{1}(0,\cdot)\equiv w_{2}(0,\cdot),\quad\forall\,F\in C_{c}^{\infty}((0,T)\times\Omega).$
@@
$-\partial_{t}w_{i}+(-\Delta)^{a}w_{i}+q_{i}w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\Omega\\ w_{i}(T,\cdot)=&f\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $f\in C_{c}^{\infty}(\Omega)$. Since ${u_{i}\in\mathrm{L}^{2}\left((0,T);H^{(a-1)(2a)}(\overline{\Omega})\right)}$ for $i=1,2$ [see theorem 8, theorem 7], integration by parts (16) yields
$\int_{\Omega}w_{i}(-\Delta)^{a}u_{i}-\int_{\Omega}u_{i}(-\Delta)^{a}w_{i}=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{u_{i}}{d^{a-1}}\frac{w_{i}}{d^{a}},\quad i=1,2,$
$\int_{\Omega}\partial_{t}u_{i}w_{i}+w_{i}\partial_{t}u_{i}=\int_{\Omega}\partial_{t}\left(u_{i}w_{i}\right)=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{u_{i}}{d^{a-1}}\frac{w_{i}}{d^{a}},\quad i=1,2.$
Integrating with respect to $t$ yields:
$\int_{\Omega}w_{i}(T,\cdot)u_{i}(T,\cdot)\,\rm{d}x=-\Gamma(a)\Gamma(a-1)\int_{0}^{T}\int_{\partial\Omega}\frac{u_{i}}{d^{a-1}}\frac{w_{i}}{d^{a}}\,\rm{d}\sigma\,\rm{d}t,\quad i=1,2,$
Substracting, we obtain
$\int_{\Omega}w_{1}(T,\cdot)u_{1}(T,\cdot)-w_{2}(T,\cdot)u_{2}(T,\cdot)\,\rm{d}x=-\Gamma(a)\Gamma(a-1)\int_{0}^{T}\int_{\partial\Omega}F\left(\frac{w_{1}}{d^{a}}-\frac{w_{2}}{d^{a}}\right)\,\rm{d}\sigma\,\rm{d}t.$
As $u_{1}(T,\cdot)\equiv u_{2}(T,\cdot)$, the above relation implies
(30) $\int_{0}^{T}\int_{\partial\Omega}F\left(\frac{w_{1}}{d^{a}}-\frac{w_{2}}{d^{a}}\right)\,\rm{d}\sigma\,\rm{d}t=0\implies\frac{w_{1}}{d^{a}}=\frac{w_{2}}{d^{a}}:=G(\text{say})$
as $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$ arbitrary. The difference $w_{1}-w_{2}$ satisfies

### v3

@@ 2. Density Results:
where $\epsilon>0$ is a fix quantity, $h\in\mathrm{L}^{2}(\Omega)$ is a given function and $)\eta_{\tilde{T}}\in C_{c}^{\infty}((0,\tilde{T}$)) such that $\eta_{\tilde{T}}\equiv 1$ in $T_{1},\tilde{T}]\subset(0,\tilde{T}$]. Furthermore $u_{F^{*}}$ and $F^{*}$ are connected by (14) with $F^{*}$ as the initial data at $T=\tilde{T}$. We intend to show the functional ${\mathcal{J}_{\epsilon,a}:\mathrm{L}^{2}(\Omega)\to\mathbb{R}\cup\{+\infty\}}$ is lower semicontinuous, strictly convex and coercive. Just like the case of theorem 2 continuity of $\mathcal{J}_{\epsilon,a}$ follows from the eigenvalue representation and (8). Here (8) provides the boundary continuity. Strict convexity follows from the strict convexity of the $\mathrm{L}^{2}$-norm and oberservability estimate Proposition 10. We only prove coercivity of the functional $\mathcal{J}_{\epsilon,a}$.
@@
    Coercivity: Let $F_{m}^{*}\in\mathrm{L}^{2}(\Omega)$ for all $m\in\mathbb{N}$ and ${\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}\to+\infty}$ as $m\to+\infty$. Consider the functions ${G_{m}^{*}=\frac{F_{m}^{*}}{\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}$. Then ${\|G_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}=1}$ for all $m\in\mathbb{N}$. Let $u_{G^{*}_{m}}$ be the solution to (14) with respect to the initial data $G_{m}^{*}$. Then ${u_{G^{*}_{m}}=\frac{u_{F^{*}_{m}}}{\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}$. We divide the functional $\mathcal{J}_{\epsilon,a}$ by $\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}$. It yields
@@
    The second scenario is ${\liminf_{m\in\mathbb{N}}\int_{0}^{\tilde{T}}\int_{\partial\Omega}\eta_{\tilde{T}}^{2}\left|\frac{u_{G_{m}^{*}}}{d^{s}}\right|^{2}\,\rm{d}x\,\rm{d}t=0}$. Thanks to the observability estimate Proposition 10, we have that ${\|u_{G_{m}^{*}}(t)\|_{\mathrm{L}^{2}(\Omega)}\to 0}$ as $m\to+\infty$, for all $t\in[0,\tilde{T}$]. Here the end points are also included because of continuity in time variable. We can represent the solution $u_{G_{m}^{*}}$ as
@@
    Hence ${\langle\phi_{n},G_{m}^{*}\rangle\to 0}$ as $m\to+\infty$. As $G_{m}^{*}$ is uniformly bounded in $\mathrm{L}^{2}$, there exists $G^{*}\in\mathrm{L}^{2}(\Omega)$ such that $G_{m}^{*}\rightharpoonup G^{*}$ weakly in $\mathrm{L}^{2}(\Omega)$. This further implies
@@
Continuity, strict convexity and coercivity helps us conclude that $\mathcal{J}_{\epsilon,a}$ attains it’s minimum for some $F_{min}^{*}\in\mathrm{L}^{2}(\Omega)$. We fix a function $G^{*}\in\mathrm{L}^{2}(\Omega)$ and let $u_{G^{*}}$ be the solution to (14) corresponding to the initial data $G^{*}$. We have
@@
$\frac{1}{2}$ $\int_{0}^{\tilde{T}}\int_{\partial\Omega}2\alpha\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t+\frac{1}{2}\int_{0}^{\tilde{T}}\int_{\partial\Omega}\alpha^{2}\eta_{\tilde{T}}^{2}\frac{(u_{G^{*}})^{2}}{d^{2a}}\,\rm{d}x\,\rm{d}t$
@@
$\int_{0}^{\tilde{T}}\int_{\partial\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t+\epsilon\frac{\int_{\Omega}F^{*}_{min}G^{*}\,\rm{d}x}{\|F_{min}^{*}\|_{\mathrm{L}^{2}(\Omega)}}-\int_{\Omega}hG^{*}\,\rm{d}x\geq 0.$
Employing Cauchy-Schwartz inequality, we obtain
$\int_{0}^{\tilde{T}}\int_{\partial\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t-\int_{\Omega}hG^{*}\,\rm{d}x\geq-\epsilon\|G^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
@@
$\int_{0}^{\tilde{T}}\int_{\partial\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t-\int_{\Omega}hG^{*}\,\rm{d}x\leq\epsilon\|G^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
@@
(26) $\left|\int_{0}^{\tilde{T}}\int_{\partial\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t-\int_{\Omega}hG^{*}\,\rm{d}x\right|\leq\epsilon\|G^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
We choose ${F=\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}}$ for the singular boundary data in (2). Let $u_{F}$ be the solutions to (2) with the singular boundary data $F$. Integration by parts (16) yields that:
@@
$\left|\int_{\Omega}\big(u_{F}(\tilde{T},\cdot)-h\big)G^{*}\,\rm{d}x\right|\leq\epsilon\|G^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
@@
$\|u_{F}(\tilde{T},\cdot)-h\|_{\mathrm{L}^{2}(\Omega)}\leq\epsilon.$
Now we move onto the proofs of the Calderón type inverse problems for the non-local heat equations. We recover the potentials by analyzing the solution at a particular time slice for (1) and by analyzing the solution on the boundary for (2).
@@ 3. Calderón type Inverse Problems:
We start with establishing the equivalency of the two Calderón type inverse problems as described in Theorem 4 and Theorem 5 respectively.
@@
$\left\{\begin{aligned} -\partial_{t}w_{i}+(-\Delta)^{a}w_{i}+q_{i}w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{N}\setminus\overline{\Omega}\\ \lim\limits_{\begin{subarray}{c}x\to y\\ y\in\partial\Omega\end{subarray}}\frac{w_{i}}{d^{a-1}}(y)=&F\qquad\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ w_{i}(T,\cdot)=&0\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right$.
where $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$. Since ${w_{i}\in\mathrm{L}^{2}\left((0,T);H^{(a-1)(2a)}(\overline{\Omega})\right)}$ for $i=1,2$ [see theorem 8, theorem 7], integration by parts (16) yields
@@
$\int_{\Omega}w_{i}\partial_{t}u_{i}+w_{i}\partial_{t}u_{i}=\int_{\Omega}\partial_{t}\left(u_{i}w_{i}\right)=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{w_{i}}{d^{a-1}}\frac{u_{i}}{d^{a}},\quad i=1,2.$
@@
$\left\{\begin{aligned} -\partial_{t}w_{i}+(-\Delta)^{a}w_{i}+q_{i}w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{N}\setminus\overline{\Omega}\\ \lim\limits_{\begin{subarray}{c}x\to y\\ y\in\partial\Omega\end{subarray}}\frac{w_{i}}{d^{a-1}}(y)=&F\qquad\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ w_{i}(T,\cdot)=&0\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right$.
where $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$. Let
$w_{1}(0,\cdot)\equiv w_{2}(0,\cdot),\quad\forall\,F\in C_{c}^{\infty}((0,T)\times\partial\Omega).$
@@
$\left\{\begin{aligned} \partial_{t}v+(-\Delta)^{a}v+(q_{1}w_{1}-q_{2}w_{2})=&0\qquad\mbox{in}\ (0,T)\times\Omega\\ v=&0\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{N}\setminus\Omega\\ v(0,\cdot)=&0\qquad\mbox{in}\ \{0\}\times\Omega.\end{aligned}\right$.
Let $V$ satisfy